2.2 MODIFIED Two forces are applied as shown to a bracket support 670 lb a) Determine the magnitude and direction of the resultant force using the triangle rule Show that a second triangle (forces in a different order) will yield the same result. 60° b) 25° 500 lb
Answer #1

a)

670 lb 60 9 300 Ib-

\underset{A}{\rightarrow}= - (500 Cos25) i + (500 Sin25) j = - 453.2 i + 211.3 j

\underset{B}{\rightarrow}= (670 Cos60) i + (670 Sin60) j = 335 i + 580.23 j

\underset{R}{\rightarrow} = ?

using triangle law

\underset{R}{\rightarrow} = \underset{A}{\rightarrow} + \underset{B}{\rightarrow}

\underset{R}{\rightarrow} = (- 453.2 i + 211.3 j) + (335 i + 580.23 j )

\underset{R}{\rightarrow} = - 118.2 i + 791.5 j

magnitude of resultant using pythagorean theorem is given as

|\underset{R}{\rightarrow}| = sqrt((- 118.2)2 + (791.5)2) = 800.3 lb

\theta = 180 - tan-1(791.5/118.2) = 98.5 deg

b)

\underset{A}{\rightarrow}= - (500 Cos25) i + (500 Sin25) j = - 453.2 i + 211.3 j

\underset{B}{\rightarrow}= (670 Cos60) i + (670 Sin60) j = 335 i + 580.23 j

\underset{R}{\rightarrow} = ?

using triangle law

\underset{R}{\rightarrow} = \underset{B}{\rightarrow} + \underset{A}{\rightarrow}

\underset{R}{\rightarrow} =  335 i + 580.23 j + - 453.2 i + 211.3 j

\underset{R}{\rightarrow} = - 118.2 i + 791.5 j

magnitude of resultant using pythagorean theorem is given as

|\underset{R}{\rightarrow}| = sqrt((- 118.2)2 + (791.5)2) = 800.3 lb

\theta = 180 - tan-1(791.5/118.2) = 98.5 deg